Problem 8 (18 points) An individual wanted to determine the relation that might exist between speed and miles per gallon of an automobile. Let Xbe the average speed of a car on the highway measured in miles per hour andlet Y represent the miles per gallon of the automobile The following data is collected 50 28 60 62 20 65 26 25 20 17
d Predict the miles per gallon of a car traveling 63 miles per hour e Predict the average speed of a car whose fuel mileage is 23 miles per gallon f Find r squared. What does this statistic tell us about between average speed and miles per gallon?
Answer #1

We now draw up the following table for using method of least squares

Sr. No.

x

x2

y

xy

1

50

2500

28

1400

2

55

3025

26

1430

3

55

3025

25

1375

4

60

3600

22

1320

5

60

3600

20

1200

6

62

3844

20

1240

7

65

4225

17

1105

8

65

4225

15

975

N = 8

∑x=472

∑x2=28044

∑y=173

∑xy=10045

∑y = aN+b∑x

173 = 8a + 472b                ……(i)

∑xy = a∑x + b ∑x2

10045 =472a + 28044b   …(ii)

Solving (i) and (ii) we get,

a = 70.39                              b = -0.826

now the line of regression of x on y is given as,

y = a + bx             i.e. y = 70.39 – 0.826x

a. To find y at x = 63

Y = 70.39 – 0.826(63) = 18.352 mils/gal

b. To find x at y = 23

23 = 70.39 – 0.826x

x = 57.372 miles /hr​​​​​​

c. To find R2

x

y

(y-y*)2

y'

(y-y’)2

50

28

40.64

29.09

55.73

55

26

19.14

24.96

11.12

55

25

11.39

24.96

11.12

60

22

0.14

20.83

0.63

60

20

2.64

20.83

0.63

62

20

2.64

19.18

5.99

65

17

21.39

16.70

24.26

65

15

43.89

16.70

24.26

Mean (y*)

21.625

∑ = 141.88

∑ = 133.73

R2133.73 = 141.88 = 0.943

The value of R2indicates that the regression line matches very closely the actual observed values. i.e it is 94.3 % accurate